OpenStudy (anonymous):

Patty and Mike are installing new floors in their house. Working alone, Mike can complete the floor in 10 hours. Patty can complete the same floor in 8 hours if working alone. How long will it take them, working together, to finish the floor? Round your answer to the nearest hundredth if necessary. The answers are a) 9.00 b) 0.23 c) 4.44 d) 18.00

OpenStudy (liam2842):

the answer is c

OpenStudy (anonymous):

Thank you I will best rspns if crrct

OpenStudy (liam2842):

\[\frac{ 10\times8 }{ 10+8 }\]

OpenStudy (liam2842):

thats the work if you need it

OpenStudy (anonymous):

Thx :)

OpenStudy (liam2842):

no problem!

OpenStudy (mathstudent55):

It will take x hours for both to complete the job working together. In 1 hour: Mike will complete 1/10 of the job Patty will complete 1/8 of the job In 1 hour, Mike completes 1/10 of x, of x/10 In 1 hour, Patty completes 1/8 of x, or x/8 When they work together, the do the entire job, or 1, so \(\dfrac{x}{10} + \dfrac{x}{8} = 1\) \(8 \times 10 \times \dfrac{x}{10} + 8 \times 10 \times\dfrac{x}{8} = 8 \times 10 \times1\) \(8x + 10x = 80\) \(18x = 80\) \(x = \dfrac{80}{18}\) \(x = 4.44\)